# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ Abstract
LASSO introduces shrinkage bias into estimated coefficients, which can adversely affect the desirable asymptotic normality and invalidate the standard inferential procedure based on the t-statistic. The desparsified LASSO has emerged as a well-known remedy for this issue. In the context of high dimensional predictive regression, the desparsified LASSO faces an additional challenge: the Stambaugh bias arising from nonstationary regressors. To restore the standard inferential procedure, we propose a novel estimator called IVX-desparsified LASSO (XDlasso). XDlasso eliminates the shrinkage bias and the Stambaugh bias simultaneously and does not require prior knowledge about the identities of nonstationary and stationary regressors. We establish the asymptotic properties of XDlasso for hypothesis testing, and our theoretical findings are supported by Monte Carlo simulations. Applying our method to real-world applications from the FRED-MD database — which includes a rich set of control variables — we investigate two important empirical questions: (i) the predictability of the U.S. stock returns based on the earnings-price ratio, and (ii) the predictability of the U.S. inflation using the unemployment rate.
Key words: Forecast, Data-rich Environment, Hypothesis testing, Shrinkage, Unit root
@@
If modern techniques would have been useless in predicting the last depression, is there any reason to believe they would do better predicting the next? The authors say that the paucity of data available… made it impractical for them to construct a more complicated and more insightful model of the economy. A richer model, Mr. [Ray] Fair concedes, might have yielded better forecasts.”
“Economic Scene; Forecasters’ Art In 1929 and Now” by Peter Passell. New York Times, August 31, 1988,
@@ 1 Introduction
Prediction is one of the fundamental tasks in statistical analysis of economic and financial data. About one century ago, quantitative analysis and forecasting services were made available to businesses and the general public (Dominguez et al.,, 1988). Irving Fisher (1867-1947) spearheaded this initiative, pioneering data-driven forecasting practices grounded in theoretical foundations. His renowned theories, namely the monetary quantity theory, inflation-deflation theory, and index theory, were supported by statistical evidence. For example, Fisher, (1925) was one of the earliest attempts to understand the source of business cycles in association with the price level (inflation), and Fisher, (1926) was the first to report “a statistical relation between unemployment and price changes”, today better known as the Phillips curve. Inference is crucial for integrating economic theory and empirical data, which has become a tradition upheld by successive generations of researchers. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in inference have uncovered new empirical patterns in prediction practices using datasets with temporal features.
@@
This paper aims at a plain quest: in a high dimensional linear predictive regression model when the number of potential regressors is larger than the sample size, how to conduct asymptotically valid inference for a regressor of particular interest. No paper has solved this question before. The challenges are twofold. Firstly, conventional predictive regressions were studied in the low-dimensional context. The notable feature of the predictive regression includes the persistence of the regressors (Stambaugh,, 1999), which upends the standard inferential theory based on the standard $t$-statistics. Secondly, to cope with the high dimensionality, one must estimate the coefficient via some regularization method, for example, when the underlying true regression model is sparse, LASSO (Tibshirani,, 1996) is the off-the-shelf method. It is well-known that the LASSO estimator is biased toward zero due to shrinkage, and follows a non-standard asymptotic distribution that is distinct from the asymptotic normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference based on the $t$-statistic, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors, and the shrinkage bias due to LASSO’s absolute-value penalty.
@@
The above diagnosis hints at a potential solution. In low-dimensional predictive regressions, Phillips and Magdalinos, (2009)’s IVX method leverages a self-generated instrument to alleviate the persistence of the regressor and thereby overcome the Stambaugh bias. In high dimensional cross sectional regressions, Zhang and Zhang, (2014)’s desparsified LASSO (Dlasso) constructs a score vector for the parameter of interest and removes the shrinkage bias via an auxiliary regression. Moreover, both methods provide estimators with asymptotic normal distributions in their respective environments.
@@
Can we alloy these two methods into one procedure to solve the inference of high dimensional predictive regression? We find the answer is no and yes. “No” is in the sense that a naive combination of them does not lead to the desirable results. The affirmative “Yes”, on the other hand, is established upon a deep understanding of the mechanism of both components and adapting them into the specific environment. This research culminates in a new IVX-Desparsified LASSO (XDlasso) estimator that is free from both biases and has an asymptotic normal distribution.
@@
With a predictor of interest in mind, the construction of XDlasso, detailed in Algorithm 1 of Section 2.3, is summarized as follows. First, a workhorse estimator is needed to level the ground for high dimensional predictive regression. When both nonstationary and stationary predictors are present in the regression, recently Mei and Shi, (2024) established the consistency of the standardized LASSO (Slasso) so this becomes a natural candidate for the workhorse. Second, the common practice of generating the instrument in IVX is insufficient — the IV must be scale-standardized to have the stochastic order aligned with all other predictors in Slasso. Finally, the standardized IV (SIV) serves as the target variable for the auxiliary Slasso regression in Dlasso to estimate the shrinkage bias. The XDlasso estimator of the parameter of interest is defined by the bias correction term added to the initial Slasso estimator, and the companion $t$-statistic is employed for statistical inference by comparing it with critical values drawn from the quantiles of the standard normal distribution.
@@
To tackle persistent regressors, the self-generated instrument of IVX in the second step is the key ingredient. This generated IV is less persistent than the unit root. This important feature enables us to decorrelate the IV and other covariates in the auxiliary Slasso regression in the third step, so that the XDlasso estimator boasts these two properties: (i) It is free from the Stambaugh bias and thus enjoys asymptotic normality; (ii) It reduces the order of shrinkage bias to make it correctable. In contrast, the ordinary Dlasso encounters a spurious auxiliary Slasso regression, failing to correct the bias arising from persistent regressors (See Section 2.4 for details). More importantly, XDlasso inference does not require a priori knowledge about the persistence of the regressor of interest and is thus free from pretesting bias. To the best of our knowledge, this is the first methodology to handle the inferential problem in high dimensional predictive regressions ($p\gg n$) with nonstationary predictors. This is also the first paper refining IVX technique for the high dimensional framework.
@@
This paper establishes asymptotic normality for our proposed XDlasso estimator and the convergence rate of its standard error. Specifically, the asymptotically normally distributed XDlasso estimator is $\sqrt{n}$-consistent for a stationary regressor while its convergence accelerates to a faster rate for a unit root regressor. Monte Carlo simulations show that XDlasso successfully removes the bias for inference on the coefficient of a nonstationary regressor, but the ordinary Dlasso does not. Our procedure is applied to the high dimensional macroeconomic FRED-MD dataset (McCracken and Ng,, 2016) with both stationary and persistent variables, studying two important macro-finance problems: financial market return predictability and the Phillips curve in macroeconomics, where certain predictors of interest in both cases are highly persistent.
@@
Literature review. This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (Bickel et al.,, 2009). It is well received and used for economic applications; Belloni et al., (2012), Shi, (2016), Caner and Kock, (2018), and Babii et al., (2022), to name a few. In recent years, the properties of LASSO estimators are studied in various topics in high dimensional time series, including nonstationary time series models (Koo et al.,, 2020; Lee et al.,, 2022; Mei and Shi,, 2024) and inference based on the heteroskedasticity and autocorrelation consistent (HAC) estimation (Babii et al.,, 2020, 2024).
@@
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, Zhang and Zhang, (2014), van de Geer et al., (2014), and Javanmard and Montanari, (2014) have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. Adamek et al., (2023) generalize the Dlasso inference to high dimensional stationary time series. We follow this line of desparsified LASSO literature considering its convenience, typically only requiring a baseline regression and an auxiliary regression. An alternative method for statistical inference is Chernozhukov et al., (2018)’s double machine learning (DML), which has been widely used in cross sectional data (Chernozhukov et al., 2022a, ; Chernozhukov et al., 2022b, ; Fan et al.,, 2024). DML adopts sample splitting or cross fitting to correct the bias in the initial estimator. In our environment with nonstationary time series, however, it is not clear how to split the sample over the time dimension, due to the time varying nature of the data generating processes.
@@
The other strand is the vast literature on predictive regressions. As highlighted by Campbell and Yogo, (2006) and Jansson and Moreira, (2006), non-standard distortion in the asymptotic distribution arises from persistent regressors. The peculiar asymptotic distributions disable the standard inferential procedures. There have been multiple proposals for valid inference, for example, the Bonferroni’s method (Campbell and Yogo,, 2006), the conditional likelihood method (Jansson and Moreira,, 2006), the linear projection method (Cai and Wang,, 2014), the weighted empirical likelihood approach (Zhu et al.,, 2014; Liu et al.,, 2019; Yang et al.,, 2021) and the implication-based inference (Xu,, 2020). Some of these methods are designed for univariate predictive regressions; it would be difficult to carry them over into the high dimensional case where the regularization method is required to cope with the issue of many parameters. On the other hand, Phillips and Magdalinos, (2009)’s IVX estimator gained its popularity by recovering the asymptotic normal distribution, enabling valid inference for mean regressions (Kostakis et al.,, 2015, 2018; Phillips and Lee,, 2013, 2016; Yang et al.,, 2020; Demetrescu et al.,, 2023), and quantile regressions (Lee,, 2016; Fan and Lee,, 2019; Cai et al.,, 2023; Liu et al.,, 2023) with multiple regressors. IVX restores asymptotic normality in low dimension by projecting the persistent regressor onto a self-generated IV, which is a more promising route to solve the problem in high dimension.
With the advent of big data, machine learning methods have spread to time series topics such as nonstationarity (Phillips and Shi,, 2021; Smeekes and Wijler,, 2021; Mei et al.,, 2024), cointegration testing (Onatski and Wang,, 2018; Zhang et al.,, 2019; Bykhovskaya and Gorin,, 2022), and structural breaks (Deshpande et al.,, 2023; Tu and Xie,, 2023).
@@
Layout. The rest of the paper is organized as follows. Section 2 introduces the high dimensional predictive regression model with a mixture of stationary and nonstationary regressors and proposes XDlasso. Section 3 establishes the theoretical results justifying the size and power of the XDlasso inference procedure. Section 4 carries out simulation studies that corroborate the theory. Section 5 applies XDlasso inference to two macro-finance empirical examples. Section 6 concludes the paper. Technical proofs are relegated to the Online Appendices.
@@
Notations. We set up the notations before formal discussions. We define $\boldsymbol{1}\left\{\cdot\right\}$ as the indicator function, and $\Delta$ as the difference operator so that $\Delta x_{t}=x_{t}-x_{t-1}$. The set of natural numbers, integers, and real numbers are denoted as $\mathbb{N}$, $\mathbb{Z}$, and $\mathbb{R}$, respectively. For some $n\in\mathbb{N}$, the integer set $\{1,2,\cdots,n\}$ is denoted as $[n]$, and the space of $n$-dimensional vectors is denoted as $\mathbb{R}^{n}.$ For $x=(x_{t})_{t\in[n]}\in\mathbb{R}^{n}$, the $L_{1}$-norm is $\left\|x\right\|_{1}=\sum_{t=1}^{n}\left|x_{t}\right|$, and the sup-norm is $\|x\|_{\infty}=\sup_{t\in[n]}|x_{t}|$. Let $0_{n}$ be an $n\times 1$ zero vector, $1_{n}$ be an $n\times 1$ vector of ones, and $I_{n}$ be the $n\times n$ identity matrix. For a generic matrix $B,$ let $B_{ij}$ be the $(i,j)$-th element, and $B^{\top}$ be its transpose. Let $\|B\|_{\infty}=\max_{i,j}|B_{ij}|$, and $\lambda_{\min}(B)$ and $\lambda_{\max}(B)$ be the minimum and maximum eigenvalues, respectively. Define $a\wedge b:=\min\left\{a,b\right\}$, and $a\vee b:=\max\left\{a,b\right\}$. An absolute constant is a positive, finite constant that is invariant with the sample size. The abbreviation “w.p.a.1” is short for “with probability approaching one”. We use $\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}$ and $\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ to denote convergence in probability and in distribution, respectively. For any time series $\{a_{t}\}_{t=1}^{n}$, we use $\bar{a}$ to denote its sample mean $n^{-1}\sum_{t=1}^{n}a_{t}$. For any time series $\{a_{t}\}$ and $\{b_{t}\}$, we say they are asymptotically uncorrelated if their sample correlation coefficient $\frac{\sum_{t=1}^{n}(a_{t}-\bar{a})(b_{t}-\bar{b})}{\sqrt{\sum_{t=1}^{n}(a_{t}-\bar{a})^{2}\sum_{t=1}^{n}(b_{t}-\bar{b})^{2}}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$ as $n\to\infty.$

### v3

# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shid

@@ Abstract
LASSO inflicts shrinkage bias on estimated coefficients, which undermines asymptotic normality and invalidates standard inferential procedures based on the t-statistic. Given cross sectional data, the desparsified LASSO has emerged as a well-known remedy for correcting the shrinkage bias. In the context of high dimensional predictive regression, the desparsified LASSO faces an additional challenge: the Stambaugh bias arising from nonstationary regressors modeled as local unit roots. To restore standard inference, we propose a novel estimator called IVX-desparsified LASSO (XDlasso). XDlasso simultaneously eliminates both shrinkage bias and Stambaugh bias and does not require prior knowledge about the identities of nonstationary and stationary regressors. We establish the asymptotic properties of XDlasso for hypothesis testing, and our theoretical findings are supported by Monte Carlo simulations. Applying our method to real-world applications from the FRED-MD database, we investigate two important empirical questions: (i) the predictability of the U.S. stock returns based on the earnings-price ratio, and (ii) the predictability of the U.S. inflation using the unemployment.
Key words: Data-rich Environment, Forecast, Hypothesis testing, LASSO, Local Unit root, Shrinkage
@@ 1 Introduction
The evaluation of economic and financial predictability has attracted widespread interest from theoretical and applied researchers for many decades. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in inference with high dimensional data have uncovered new empirical patterns in predictive practices using large-scale datasets with temporal features.
@@
This paper aims at a plain quest: in a high dimensional linear predictive regression model, where the number of potential regressors is larger than the sample size, how can one conduct valid inference for a regressor of primary interest? No research has solved this question before. The challenges are twofold. First, predictive regressions were mainly studied in the low-dimensional context. A defining feature of predictive regression theory lies in persistent regressors (stambaugh1999predictive), which upend standard inference based on the standard $t$-statistics. Second, one must estimate the coefficient via some regularization methods to cope with high dimensionality. For example, when the underlying true regression model is sparse, LASSO (tibshirani1996regression) is the off-the-shelf method. It is well known that the LASSO estimator is biased toward zero due to absolute-value shrinkage and exhibits a nonstandard asymptotic distribution that is distinct from the normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors and the shrinkage bias caused by the LASSO penalty.
@@
The above diagnosis hints at a plausible solution path. In low-dimensional predictive regressions, phillips2009econometric’s IVX method leverages a self-generated instrument to alleviate the regressor persistence, thereby overcoming the Stambaugh bias. In high dimensional cross sectional regressions, zhang2014confidence’s desparsified LASSO (Dlasso) constructs a score vector for the parameter of interest and removes the shrinkage bias via an auxiliary regression. Each method provides an asymptotic normal estimator in its respective environment.
@@
Can we combine these two methods into a single procedure to address inference in high dimensional predictive regressions? We find that the answer is both no and yes. “No” is in the sense that a naive combination of the two does not lead to desirable results. “Yes”, on the other hand, is established upon a deep understanding of the mechanisms of both components and their adaptation to the context. This research culminates in a new IVX-Desparsified LASSO (XDlasso) estimator that is free from both biases and has an asymptotic normal distribution.
With a predictor of interest in mind, the construction of XDlasso, detailed in Algorithm 1 in Section 2.3, is summarized as follows. First, a workhorse estimator is needed to lay the groundwork for a high dimensional predictive regression. When both nonstationary and stationary predictors are present in the regression, mei2022lasso has recently established the consistency of the standardized LASSO (Slasso), making it a natural candidate for the workhorse estimator. Beyond the pure unit roots considered in MS24, we extend the characterization of nonstationary regressors by generalizing our framework to allow for local unit root (LUR) processes. The convergence rates of Slasso for LURs align with those for pure unit roots. However, the technical proofs for LURs are more involved than those for pure unit roots. We address the complexity arising from high dimensional predictors with both LURs and stationary regressors, and derive the convergence rates of the initial Slasso estimator.
@@
Second, the common practice of generating the instrument in IVX is insufficient — the IV must be scale-standardized to have the stochastic order aligned with all other predictors in Slasso. The standardized IV serves as the target variable for the auxiliary Slasso regression in Dlasso to estimate the shrinkage bias. The XDlasso estimator of the parameter of interest is defined as the initial Slasso estimator plus the bias-correction term, and the companion $t$-statistic is employed for statistical inference by comparing it to critical values from the standard normal distribution.
We further establish the asymptotic normality of our proposed XDlasso estimator and the convergence rate of its standard error. Specifically, the XDlasso estimator is $\sqrt{n}$-consistent for a stationary regressor while its convergence accelerates for an LUR regressor. Moreover, to conduct simultaneous inference for multiple parameters of interest, we develop a Wald statistic with an asymptotic $\chi^{2}$ distribution based on XDlasso. This Wald test is valid even when the parameters involve both stationary and nonstationary regressors.
@@
To tackle persistent regressors, the self-generated IVX instrument in the second step is the key ingredient. The generated IV is less persistent than the nonstationary regressors modeled as LURs. This important feature enables us to decorrelate the IV from other covariates in the auxiliary LASSO regression, so that the resulting XDlasso estimator possesses these two properties: (i) It is free from the Stambaugh bias and thus enjoys asymptotic normality; (ii) It reduces the order of shrinkage bias to make it correctable. In contrast, the ordinary Dlasso encounters a spurious auxiliary regression, failing to correct the bias arising from persistent regressors (see Section 2.4 for details). More importantly, XDlasso inference does not require a priori knowledge of the persistence of the regressor of interest and is thus immune to pretesting bias. To the best of our knowledge, this is the first methodology to handle the inferential problem in high dimensional predictive regressions ($p\gg n$) with nonstationary predictors. This is also the first paper that extends the IVX technique into the high dimensional framework.
@@
Monte Carlo simulations show that XDlasso successfully removes the bias for inference on the coefficient of a nonstationary regressor, but the ordinary Dlasso fails to do so. Our procedure is applied to the high dimensional macroeconomic FRED-MD dataset (mccracken2016fred) with both stationary and persistent variables, to study two important macro-finance problems: financial market return predictability and the Phillips curve in macroeconomics.
@@
Literature review. With the advent of big data, machine learning methods have spread to time series topics such as nonstationarity (phillips2021boosting; smeekes2021automated; mei2024boosted), cointegration testing (onatski2018alternative; zhang2019identifying; bykhovskaya2022cointegration), and structural breaks (deshpande2023online; tu2023penetrating). This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (bickel2009simultaneous). It is well received and used for economic applications; see belloni2012sparse, shi2016estimation, caner2018asymptotically, and babii2022machine, to name a few. In recent years, the properties of LASSO are studied in various topics in high dimensional time series, including nonstationary time series models (koo2020high; lee2022lasso) and inference based on the heteroskedasticity and autocorrelation consistent (HAC) estimation (babii2021highdimensionalgrangercausalitytests; babii2024high). None of these works has considered hypothesis testing problem for high dimensional predictive regressions with both LURs and stationary regressors.
@@
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, zhang2014confidence, van2014asymptotically, and javanmard2014confidence have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. adamek2022lasso generalize the Dlasso inference to high dimensional stationary time series. We follow this line of desparsified LASSO literature thanks to its convenience, which requires a baseline regression and an auxiliary regression only. On the other hand, chernozhukov2018’s double machine learning (DML) is a more general theoretical framework of debiased inference, widely used in cross sectional data where sample-splitting is readily implementable. However, none of the aforementioned works has devised any inferential procedure for high dimensional nonstationary time series. hecq2023inference apply a post-double selection procedure to test the Granger causality in high dimensional nonstationary vector autoregressive models with cointegrated data. In contrast, our procedure relies on desparsified LASSO without variable selection.
@@
The other strand is the vast literature on predictive regressions. As highlighted by campbell2006efficient and jansson2006optimal, non-standard distortion in the asymptotic distribution arises from persistent regressors. The peculiar asymptotic distributions invalidate the standard inferential procedures. There have been multiple proposals for valid inference, for example, the Bonferroni method (campbell2006efficient), the conditional likelihood method (jansson2006optimal), the linear projection method (cai2014testing), the weighted empirical likelihood approach (zhu2014predictive; liu2019unified; yang2021unified), and the implication-based inference (xu2020testing). Some of these methods are designed for univariate predictive regressions; it would be difficult to extend them to the high dimensional case, where regularization is required to handle many parameters. On the other hand, phillips2009econometric’s IVX estimator gained its popularity by recovering asymptotic normality, enabling valid inference for mean regressions (kostakis2015robust; kostakis2018taking; phillips2013predictive; phillips2016robust; yang2020testing; demetrescu2023extensions) and quantile regressions (lee2016predictive; fan2019predictive; cai2023new; liu2023unified) with low dimensional regressors. IVX recovers asymptotic normality by projecting the persistent regressor onto a self-generated IV.
@@
Layout. The rest of the paper is organized as follows. Section 2 introduces the high dimensional predictive regression model with a mixture of stationary and nonstationary regressors and proposes XDlasso. Section 3 establishes the theoretical results, justifying the size and power of the XDlasso inference procedure. Section 4 carries out simulation studies that corroborate the theory. Section 5 applies XDlasso inference to two macro-finance empirical examples. Technical proofs are relegated to the Online Appendices.
@@
Notations. We set up the notation before the formal discussion. We define $\boldsymbol{1}\left\{\cdot\right\}$ as the indicator function, and $\Delta$ as the difference operator so that $\Delta x_{t}=x_{t}-x_{t-1}$. The set of natural numbers, integers, and real numbers are denoted as $\mathbb{N}$, $\mathbb{Z}$, and $\mathbb{R}$, respectively. For some $n\in\mathbb{N}$, the integer set $\{1,2,\cdots,n\}$ is denoted as $[n]$, and the space of $n$-dimensional vectors is denoted as $\mathbb{R}^{n}.$ For $x=(x_{t})_{t\in[n]}\in\mathbb{R}^{n}$, the $L_{1}$-norm is $\left\|x\right\|_{1}=\sum_{t=1}^{n}\left|x_{t}\right|$, and the sup-norm is $\|x\|_{\infty}=\sup_{t\in[n]}|x_{t}|$. Let $0_{n}$ be an $n\times 1$ zero vector, $1_{n}$ be an $n\times 1$ vector of ones, and $I_{n}$ be the $n\times n$ identity matrix. For a generic matrix $B,$ let $B_{ij}$ be the $(i,j)$-th element, and $B^{\top}$ be its transpose. Let $\|B\|_{\infty}=\max_{i,j}|B_{ij}|$, and $\lambda_{\min}(B)$ and $\lambda_{\max}(B)$ be the minimum and maximum eigenvalues, respectively. Define $a\wedge b:=\min\left\{a,b\right\}$, and $a\vee b:=\max\left\{a,b\right\}$. An absolute constant is a positive, finite constant that is invariant with the sample size. The abbreviation “w.p.a.1” is short for “with probability approaching one”. We use $\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}$ and $\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ to denote convergence in probability and in distribution, respectively. For any time series $\{a_{t}\}_{t=1}^{n}$, we use $\bar{a}$ to denote its sample mean $n^{-1}\sum_{t=1}^{n}a_{t}$. For any time series $\{a_{t}\}$ and $\{b_{t}\}$, we say they are asymptotically uncorrelated if their sample correlation coefficient $\frac{\sum_{t=1}^{n}(a_{t}-\bar{a})(b_{t}-\bar{b})}{\sqrt{\sum_{t=1}^{n}(a_{t}-\bar{a})^{2}\sum_{t=1}^{n}(b_{t}-\bar{b})^{2}}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$ as $n\to\infty.$